# Uncovering the Mechanism of Chiral Three-Nucleon Force in Driving Spin-Orbit Splitting

### v1

@@ B.0.2 General expression of irreducible-tensor decomposition
$\bm{a}\cdot\bm{b}$ $=-\sqrt{3}\left[a_{1}\otimes b_{1}\right]_{00},$ (12)
$\left(\bm{a}\times\bm{b}\right)_{\mu}$ $=-i\sqrt{2}\left[a_{1}\otimes b_{1}\right]_{1\mu}\quad(\mu=0,\pm 1),$ (13)
where $\bm{a}$ and $\bm{b}$ are arbitrary three-dimensional vectors, which correspond to irreducible tensors of rank 1, $a_{1m}$ and $b_{1m}$, respectively. In general, the tensor product of arbitrary irreducible tensors $\mathcal{M}_{lm}$ and $\mathcal{N}_{l^{\prime}m^{\prime}}$ is defined by
$\left[\mathcal{M}_{l}\otimes\mathcal{N}_{l^{\prime}}\right]_{\lambda\mu}=\sum_{mm^{\prime}}\left(lml^{\prime}m^{\prime}|\lambda\mu\right)\mathcal{M}_{lm}\mathcal{M}_{l^{\prime}m^{\prime}},$ (14)
with the Clebsch-Gordan coefficient (....|..)(\,.\,.\,.\,.\,|\,.\,.\,)( . . . . | . . ). Furthermore, assuming $\bm{a}$ depends on neither spin variables nor differential operators, it can be represented in terms of the spherical harmonics $Y_{1m}$ [75]:
$a_{1m}=\sqrt{\frac{4\pi}{3}}aY_{1m}(\hat{\bm{a}}),$ (15)
@@
$w^{(\alpha)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\bm{q}_{i},\bm{q}_{j}\right)$
$\quad=w_{\mathrm{pro}}^{(\alpha)}(q_{i},q_{j})\sum_{\lambda}\mathcal{O}_{\lambda}^{(\alpha)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right),$ (16)
$\mathcal{O}_{\lambda}^{(\alpha)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$
$\quad=A_{\lambda}\left[\mathcal{M}_{\lambda}^{(\alpha)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k}\right)\otimes\mathcal{N}_{\lambda}^{(\alpha)}\!\left(\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)\right]_{00}.$ (17)
@@ B.0.3 Contact term
$w^{(\mathrm{ct})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\bm{q}_{i},\bm{q}_{j}\right)=\frac{c_{E}}{2f_{\pi}^{4}\Lambda_{\chi}}.$ (18)
@@
B.0.4 One-pion exchange plus contact term
@@
$w^{(1\pi)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\bm{q}_{i},\bm{q}_{j}\right)$ $=-\frac{g_{A}c_{D}}{8f_{\pi}^{4}\Lambda_{\chi}}\frac{q_{j}^{2}}{q_{j}^{2}+m_{\pi}^{2}}\sum_{\lambda=0,2}\mathcal{O}_{\lambda}^{(1\pi)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right),$ (19)
$\mathcal{O}_{\lambda}^{(1\pi)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$ $=\sqrt{4\pi}\left(1010|\lambda 0\right)\left[\left[\sigma_{1}(i)\otimes\sigma_{1}(j)\right]_{\lambda}\otimes Y_{\lambda}\!\left(\hat{\bm{q}}_{j}\right)\right]_{00}.$ (20)
@@
$\mathcal{O}_{\lambda}^{(1\pi)}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$ =$=\begin{dcases}\frac{1}{3}\bm{\sigma}_{i}\cdot\bm{\sigma}_{j}\quad&(\lambda=0),\\ \frac{1}{3}\mathcal{S}_{ij}(\hat{\bm{q}}_{j})\quad&(\lambda=2),\end{dcases}$ (21)
@@
$\mathcal{S}_{12}(\hat{\bm{q}})=3\left(\bm{\sigma}_{1}\cdot\hat{\bm{q}}\right)\left(\bm{\sigma}_{2}\cdot\hat{\bm{q}}\right)-\bm{\sigma}_{1}\cdot\bm{\sigma}_{2}.$ (22)
B.0.5 Two-pion exchange $c_{1}$ term
@@
$w^{(c_{1})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\bm{q}_{i},\bm{q}_{j}\right)$ $=-\frac{g_{A}^{2}c_{1}m_{\pi}^{2}}{2f_{\pi}^{4}}\frac{q_{i}q_{j}}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)}\sum_{\lambda=0}^{2}\mathcal{O}_{\lambda}^{(c_{1})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right),$ (23)
$\mathcal{O}_{\lambda}^{(c_{1})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$ $=\frac{4\pi}{3}\hat{\lambda}\left[\left[\sigma_{1}(i)\otimes\sigma_{1}(j)\right]_{\lambda}\otimes\left[Y_{1}\!\left(\hat{\bm{q}}_{i}\right)\otimes Y_{1}\!\left(\hat{\bm{q}}_{j}\right)\right]_{\lambda}\right]_{00}.$ (24)
@@
$\mathcal{O}_{\lambda}^{(c_{1})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$ =$=\begin{dcases}\frac{1}{3}\left(\bm{\sigma}_{i}\cdot\bm{\sigma}_{j}\right)\left(\hat{\bm{q}}_{i}\cdot\hat{\bm{q}}_{j}\right)\quad&(\lambda=0),\\ \frac{1}{2}\left(\bm{\sigma}_{i}\times\bm{\sigma}_{j}\right)\cdot\left(\hat{\bm{q}}_{i}\times\hat{\bm{q}}_{j}\right)\quad&(\lambda=1),\\ \frac{1}{3}\mathcal{T}_{ij}\left(\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)\quad&(\lambda=2).\end{dcases}$ (25)
@@
$\mathcal{T}_{12}(\hat{\bm{q}},\hat{\bm{q}}^{\prime})$ $=\frac{3}{2}[(\bm{\sigma}_{1}\cdot\hat{\bm{q}})(\bm{\sigma}_{2}\cdot\hat{\bm{q}}^{\prime})+(\bm{\sigma}_{2}\cdot\hat{\bm{q}})(\bm{\sigma}_{1}\cdot\hat{\bm{q}}^{\prime})]-(\bm{\sigma}_{1}\cdot\bm{\sigma}_{2})(\hat{\bm{q}}\cdot\hat{\bm{q}}^{\prime}).$ (26)
B.0.6 Two-pion exchange $c_{3}$ term
The $c_{3}$ potential $v_{3N}^{(c_{3})}$ given by Eq. (7) differs from $v_{3N}^{(c_{1})}$ only in the prefactor and the presence of $\bm{q}_{i}\cdot\bm{q}_{j}$ responsible for the $p$-wave pion-nucleon scattering. Therefore, the structure of $w^{(c_{3})}$ is similar to that of $w^{(c_{1})}$:
$w_{3N}^{(c_{3})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\bm{q}_{i},\bm{q}_{j}\right)$ $=\frac{g_{A}^{2}c_{3}}{4f_{\pi}^{4}}\frac{q_{i}^{2}q_{j}^{2}}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)}\sum_{\lambda=0}^{2}\mathcal{O}_{\lambda}^{(c_{3})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right),$ (27)
$\mathcal{O}_{\lambda}^{(c_{3})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$ $=4\pi(-)^{\lambda+1}\hat{\lambda}\sum_{l_{i}l_{j}}\left(1010|l_{i}0\right)\left(1010|l_{j}0\right)\begin{Bmatrix}l_{i}&l_{j}&\lambda\\ 1&1&1\end{Bmatrix}$
×$\times\left[\left[\sigma_{1}(i)\otimes\sigma_{1}(j)\right]_{\lambda}\otimes\left[Y_{l_{i}}\!\left(\hat{\bm{q}}_{i}\right)\otimes Y_{l_{j}}\!\left(\hat{\bm{q}}_{j}\right)\right]_{\lambda}\right]_{00}.$ (28)
@@ B.0.6 Two-pion exchange $c_{3}$ term
$\mathcal{O}_{\lambda}^{(c_{3})}\!\left(\bm{\sigma}_{i},\bm{\sigma}_{j},\bm{\sigma}_{k},\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)$ =$=\begin{dcases}\frac{1}{3}\left(\bm{\sigma}_{i}\cdot\bm{\sigma}_{j}\right)\left(\hat{\bm{q}}_{i}\cdot\hat{\bm{q}}_{j}\right)^{2}&(\lambda=0),\\ \frac{1}{2}\left(\bm{\sigma}_{i}\times\bm{\sigma}_{j}\right)\cdot\left(\hat{\bm{q}}_{i}\times\hat{\bm{q}}_{j}\right)\left(\hat{\bm{q}}_{i}\cdot\hat{\bm{q}}_{j}\right)&(\lambda=1),\\ \frac{1}{3}\left(\hat{\bm{q}}_{i}\cdot\hat{\bm{q}}_{j}\right)\mathcal{T}_{ij}\left(\hat{\bm{q}}_{i},\hat{\bm{q}}_{j}\right)&(\lambda=2).\\ \end{dcases}$ (29)
B.0.7 Two-pion exchange $c_{4}$ term

### v2

@@ A.2 General expression of irreducible-tensor decomposition
$\boldsymbol{a}\cdot\boldsymbol{b}$ $=-\sqrt{3}\left[a_{1}\otimes b_{1}\right]_{00},$ (12)
$\left(\boldsymbol{a}\times\boldsymbol{b}\right)_{\mu}$ $=-i\sqrt{2}\left[a_{1}\otimes b_{1}\right]_{1\mu}\quad(\mu=0,\pm 1),$ (13)
where $\boldsymbol{a}$ and $\boldsymbol{b}$ are arbitrary three-dimensional vectors, which correspond to irreducible tensors of rank 1, $a_{1m}$ and $b_{1m}$, respectively. In general, the tensor product of arbitrary irreducible tensors $\mathcal{M}_{lm}$ and $\mathcal{N}_{l^{\prime}m^{\prime}}$ is defined by
$\left[\mathcal{M}_{l}\otimes\mathcal{N}_{l^{\prime}}\right]_{\lambda\mu}=\sum_{mm^{\prime}}\left(lml^{\prime}m^{\prime}|\lambda\mu\right)\mathcal{M}_{lm}\mathcal{N}_{l^{\prime}m^{\prime}},$ (14)
with the Clebsch-Gordan coefficient (....|..)(\,.\,.\,.\,.\,|\,.\,.\,)( . . . . | . . ). Furthermore, assuming $\boldsymbol{a}$ depends on neither spin variables nor differential operators, it can be represented in terms of the spherical harmonics $Y_{1m}$ [81]:
$a_{1m}=\sqrt{\frac{4\pi}{3}}aY_{1m}(\hat{\boldsymbol{a}}),$ (15)
@@
$w^{(\alpha)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\boldsymbol{q}_{i},\boldsymbol{q}_{j}\right)$
$\quad=w_{\mathrm{pro}}^{(\alpha)}(q_{i},q_{j})\sum_{\lambda}\mathcal{O}_{\lambda}^{(\alpha)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right),$ (16)
$\mathcal{O}_{\lambda}^{(\alpha)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$
$\quad=A_{\lambda}\left[\mathcal{M}_{\lambda}^{(\alpha)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k}\right)\otimes\mathcal{N}_{\lambda}^{(\alpha)}\!\left(\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)\right]_{00}.$ (17)
@@ A.2.1 Contact term
$w^{(\mathrm{ct})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\boldsymbol{q}_{i},\boldsymbol{q}_{j}\right)=\frac{c_{E}}{2f_{\pi}^{4}\Lambda_{\chi}}.$ (18)
@@
A.3 One-pion exchange plus contact term
@@
$w^{(1\pi)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\boldsymbol{q}_{i},\boldsymbol{q}_{j}\right)$
$\quad=-\frac{g_{A}c_{D}}{8f_{\pi}^{4}\Lambda_{\chi}}\frac{q_{j}^{2}}{q_{j}^{2}+m_{\pi}^{2}}\sum_{\lambda=0,2}\mathcal{O}_{\lambda}^{(1\pi)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right),$ (19)
$\mathcal{O}_{\lambda}^{(1\pi)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$
$\quad=\sqrt{4\pi}\left(1010|\lambda 0\right)\left[\left[\sigma_{1}(i)\otimes\sigma_{1}(j)\right]_{\lambda}\otimes Y_{\lambda}\!\left(\hat{\boldsymbol{q}}_{j}\right)\right]_{00}.$ (20)
@@
$\mathcal{O}_{\lambda}^{(1\pi)}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$ =$=\begin{dcases}\frac{1}{3}\boldsymbol{\sigma}_{i}\cdot\boldsymbol{\sigma}_{j}\quad&(\lambda=0),\\ \frac{1}{3}\mathcal{S}_{ij}(\hat{\boldsymbol{q}}_{j})\quad&(\lambda=2),\end{dcases}$ (21)
@@
$\mathcal{S}_{12}(\hat{\boldsymbol{q}})=3\left(\boldsymbol{\sigma}_{1}\cdot\hat{\boldsymbol{q}}\right)\left(\boldsymbol{\sigma}_{2}\cdot\hat{\boldsymbol{q}}\right)-\boldsymbol{\sigma}_{1}\cdot\boldsymbol{\sigma}_{2}.$ (22)
A.4 Two-pion exchange $c_{1}$ term
@@
$w^{(c_{1})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\boldsymbol{q}_{i},\boldsymbol{q}_{j}\right)$ $=-\frac{g_{A}^{2}c_{1}m_{\pi}^{2}}{2f_{\pi}^{4}}\frac{q_{i}q_{j}}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)}$
×$\times\sum_{\lambda=0}^{2}\mathcal{O}_{\lambda}^{(c_{1})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right),$ (23)
$\mathcal{O}_{\lambda}^{(c_{1})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$
$\quad=\frac{4\pi}{3}\hat{\lambda}\left[\left[\sigma_{1}(i)\otimes\sigma_{1}(j)\right]_{\lambda}\otimes\left[Y_{1}\!\left(\hat{\boldsymbol{q}}_{i}\right)\otimes Y_{1}\!\left(\hat{\boldsymbol{q}}_{j}\right)\right]_{\lambda}\right]_{00}.$ (24)
@@
$\mathcal{O}_{\lambda}^{(c_{1})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$ =$=\begin{dcases}\frac{1}{3}\left(\boldsymbol{\sigma}_{i}\cdot\boldsymbol{\sigma}_{j}\right)\left(\hat{\boldsymbol{q}}_{i}\cdot\hat{\boldsymbol{q}}_{j}\right)\quad&(\lambda=0),\\ \frac{1}{2}\left(\boldsymbol{\sigma}_{i}\times\boldsymbol{\sigma}_{j}\right)\cdot\left(\hat{\boldsymbol{q}}_{i}\times\hat{\boldsymbol{q}}_{j}\right)\quad&(\lambda=1),\\ \frac{1}{3}\mathcal{T}_{ij}\left(\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)\quad&(\lambda=2).\end{dcases}$ (25)
@@
$\mathcal{T}_{12}(\hat{\boldsymbol{q}},\hat{\boldsymbol{q}}^{\prime})$ $=\frac{3}{2}[(\boldsymbol{\sigma}_{1}\cdot\hat{\boldsymbol{q}})(\boldsymbol{\sigma}_{2}\cdot\hat{\boldsymbol{q}}^{\prime})+(\boldsymbol{\sigma}_{2}\cdot\hat{\boldsymbol{q}})(\boldsymbol{\sigma}_{1}\cdot\hat{\boldsymbol{q}}^{\prime})]$
$-(\boldsymbol{\sigma}_{1}\cdot\boldsymbol{\sigma}_{2})(\hat{\boldsymbol{q}}\cdot\hat{\boldsymbol{q}}^{\prime}).$ (26)
A.5 Two-pion exchange $c_{3}$ term
The $c_{3}$ potential $v_{3N}^{(c_{3})}$ given by Eq. (7) differs from $v_{3N}^{(c_{1})}$ only in the prefactor and the presence of $\boldsymbol{q}_{i}\cdot\boldsymbol{q}_{j}$ responsible for the $p$-wave pion-nucleon scattering. Therefore, the structure of $w^{(c_{3})}$ is similar to that of $w^{(c_{1})}$:
$w_{3N}^{(c_{3})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\boldsymbol{q}_{i},\boldsymbol{q}_{j}\right)$ $=\frac{g_{A}^{2}c_{3}}{4f_{\pi}^{4}}\frac{q_{i}^{2}q_{j}^{2}}{\left(q_{i}^{2}+m_{\pi}^{2}\right)\left(q_{j}^{2}+m_{\pi}^{2}\right)}$
×$\times\sum_{\lambda=0}^{2}\mathcal{O}_{\lambda}^{(c_{3})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right),$ (27)
$\mathcal{O}_{\lambda}^{(c_{3})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$
$\quad=4\pi(-)^{\lambda+1}\hat{\lambda}\sum_{l_{i}l_{j}}\left(1010|l_{i}0\right)\left(1010|l_{j}0\right)\begin{Bmatrix}l_{i}&l_{j}&\lambda\\ 1&1&1\end{Bmatrix}$
×$\quad\times\left[\left[\sigma_{1}(i)\otimes\sigma_{1}(j)\right]_{\lambda}\otimes\left[Y_{l_{i}}\!\left(\hat{\boldsymbol{q}}_{i}\right)\otimes Y_{l_{j}}\!\left(\hat{\boldsymbol{q}}_{j}\right)\right]_{\lambda}\right]_{00}.$ (28)
@@ A.5 Two-pion exchange $c_{3}$ term
$\mathcal{O}_{\lambda}^{(c_{3})}\!\left(\boldsymbol{\sigma}_{i},\boldsymbol{\sigma}_{j},\boldsymbol{\sigma}_{k},\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)$
=$\quad=\begin{dcases}\frac{1}{3}\left(\boldsymbol{\sigma}_{i}\cdot\boldsymbol{\sigma}_{j}\right)\left(\hat{\boldsymbol{q}}_{i}\cdot\hat{\boldsymbol{q}}_{j}\right)^{2}&(\lambda=0),\\ \frac{1}{2}\left(\boldsymbol{\sigma}_{i}\times\boldsymbol{\sigma}_{j}\right)\cdot\left(\hat{\boldsymbol{q}}_{i}\times\hat{\boldsymbol{q}}_{j}\right)\left(\hat{\boldsymbol{q}}_{i}\cdot\hat{\boldsymbol{q}}_{j}\right)&(\lambda=1),\\ \frac{1}{3}\left(\hat{\boldsymbol{q}}_{i}\cdot\hat{\boldsymbol{q}}_{j}\right)\mathcal{T}_{ij}\left(\hat{\boldsymbol{q}}_{i},\hat{\boldsymbol{q}}_{j}\right)&(\lambda=2).\\ \end{dcases}$ (29)
A.6 Two-pion exchange $c_{4}$ term